Following vigorous exercise, the body temperature of a 60.0 kg person is 39.9°C . At what rate in watts must the person transfer thermal energy to reduce the the body temperature to 37.0°C in 28.0 min, assuming the body continues to produce energy at the rate of 150 W? (1 watt = 1 joule/second or 1 W = 1 J/s). W
Answer #1

Q = mc dT

Q = 60 x 3500 x (39.9 - 37) = 609000 J

Power, P = 609000/28 x 60 = 362.5 W

Total power, Pt = 362.5 + 150 = 512.5 W

Comment in case any doubt please rate my answer ...